# Digi2Real: Bridging the Realism Gap in Synthetic Data Face Recognition via Foundation Models

### v2

@@ 5 Discussions
Limitations While this approach successfully generates high-quality images, there are several important spects to note. Firstly, the Arc2Face training pipeline utilizes images from WebFace42M, FFHQ, and CelebA-HQ, indicating that the Arc2Face model requires supervision from a large number of samples to generate identity-consistent images. Moreover, their method depends on a high-quality face recognition model, also trained on the WebFace dataset, which is a common limitation among many synthetic datasets. Additionally, as shown in the figures, the generated images do not always appear very realistic.
@@ 6 Conclusion
In this work, we show that transferring realism, to images generated from a graphics pipeline can greatly increase face recognition performance. The usage of the graphics pipeline makes it possible to generate virtually a large number of identities with user-defined poses, expressions, accessories, lighting, ethnicities, and so on offering fine-grained control over the generation process. This makes it possible to generate balanced datasets easily. Now, the second step in the process is to make the generated images look photorealistic, our approach can be used to make them look realistic to boost this process. Essentially, the hybrid approach we proposed combines the effectiveness of both synthetic data generation paradigms and boosts the face recognition performance further. We also note that there needs to be more innovation in transferring realism with a minimal number of real data usage. We hope that the success of the approach we demonstrated will encourage researchers to focus more on hybrid approaches to synthetic data generation.
@@ Acknowledgment
This research is based on work conducted in the SAFER project and supported by the Hasler Foundation’s Responsible AI program

### v4

@@ 5 Discussions
Limitations While this approach successfully generates high-quality images, there are several important aspects to note. Firstly, the Arc2Face training pipeline utilizes images from WebFace42M, FFHQ, and CelebA-HQ, indicating that the Arc2Face model requires supervision from a large number of samples to generate identity-consistent images. Moreover, their method depends on a high-quality face recognition model, also trained on the WebFace dataset, which is a common limitation among many synthetic datasets. Additionally, as shown in the figures, the generated images do not always appear very realistic.
@@ 6 Conclusion
In this work, we demonstrate that adding realism to images generated from a graphics pipeline can greatly enhance face recognition performance. The use of the graphics pipeline allows for the creation of a large number of identities with user-defined poses, expressions, accessories, lighting, ethnicities, etc., providing fine-grained control over the generation process. This capability facilitates the easy generation of balanced datasets. The next step is to make the generated images appear photorealistic, and our approach can be employed to achieve this realism. Essentially, our proposed hybrid approach combines the effectiveness of both graphics rendering and data-driven synthetic data generation paradigms, further boosting face recognition performance. We also note the need for more innovation in transferring realism with minimal use of real data. We hope the performance improvement with our approach will encourage researchers to focus more on hybrid approaches to synthetic data generation.
@@ Acknowledgment
This research is based on work conducted in the SAFER project and supported by the Hasler Foundation’s Responsible AI program.